OpenStudy (anonymous):

Jesse received two offers to join a music program. If he is enrolled in the program for 4 months, which program should he choose? First Program: 6 free songs for joining and 1 free song every month Second Program: 2 free songs for joining and 2 free songs per month for the first 3 months only A. Not enough information B. Either-He will get the same number of free songs from both C. The second program D. The first program

OpenStudy (anonymous):

D

OpenStudy (anonymous):

D

OpenStudy (dumbcow):

First program: 6+4 = 10 Second program: 2+(2*3) = 2+6 = 8